OpenStudy (e28195):

Select the x-intercepts for f(x) = sin (x - Π/2) from 0 to 2Π. a. 3Π/2 b. Π/2 c. 0 d. 2Π Explanation would be sweet thanks

OpenStudy (danjs):

you know when y=sin(x) is zero?

OpenStudy (danjs):

This one like the form f(x) = sin(x - a) shifts the graph by an amount 'a'

OpenStudy (danjs):

if y = sin(x) intercepts the x-axis for x=0, x=pi, x=2pi then y=sin(x - pi/2) shifts the graph pi/2 units to the right. All the intercepts are pi/2 more now, x=pi/2, x=3pi/2

OpenStudy (e28195):

thanks!